GLOBAL FIT OF THE HIGH RESOLUTION INFRARED SPECTRUM OF D_{2}S

\underline{O. N. ULENIKOV}, \underline{E. S. BEKHTEREVA}, Physical Chemistry, ETH-Zürich, CH-8093; G. A. ONOPENKO, O. V. GROMOVA, Laboratory of Molecular Spectroscopy, Physics Department, Tomsk State University, Tomsk, 634050, Russia; A.-W. LIU, L. WAN, L.-Y. HAO, S.-M. HU, Hefei National Laboratory for Physical Sciences at Microscale, University of Science and Technology of China, Hefei, 230026, China; \underline{J.-M. FLAUD}, LISA-UMR7583 CNRS, Universite Paris 12 et 7, 61 Av. du General de Gaulle, 94010, Creteil cedex, France.

High resolution Fourier transform spectra of the D_{2}S molecule in the regions of polyads of interacting vibrational states v = 3/2, 2, 5/2, 3 and 7/2 (v = v_{1} + v_{2}/2 + v_{3}) were recorded for the first time with a Bruker IFS 120 Fourier transform interferometer and analysed. A global fit of all currently available rotation-vibration energies has been made for 22 vibrational states of the D_{2}S molecule. The resulting set of 231 parameters reproduces all the initial experimental data (about 3670 vibration-rotation energies which correspond to more than 9700 ro-vibrational transitions with J^{max.} = 25) with accuracies close to the experimental uncertainties.